The diagonal of a rectangle forms a 30 degree with one side of the rectangle. If the diagonal has measure 10, what is the area of the rectangle?
Answer by edjones(8007)

The diagonal of the rectangle is the hypotenuse of a 30-60-90 degree triangle.
h=10
sin(30)=opp/10
opp=sin(30)*10=5
.
cos(30)=adj/10
adj=cos(30)*10=8.66...
.
A=5*8.66...=43.3013 sq units
.
Ed